# Federated Unsupervised Domain Generalization using Global and Local Alignment of Gradients

### v1

@@ 5 Results
Performance. We report the accuracy rates of FedGaLA and baseline models on the four datasets. As shown in Table 1, FedGaLA consistently outperforms all baselines across all four datasets, with the exception of ‘Art-painting’ domain in Office-Home, for the 10% data regime. When 30% of the data are used, our method still generally outperforms the baseline models, although for some of the domains, the baseline solutions produce slightly better results. This observation is expected given that with the introduction of more domain-specific training data, the need for domain generalization declines, and thus, methods that are not explicitly designed for domain generalization can produce competitive results. Across the four datasets, we observe that among the baseline models, FedSimCLR achieves better results compared to FedSimSiam, FedBYOL, and FedMoCo. This finding is consistent with [45] where it was demonstrated that SimCLR provides a better foundation for domain generalization versus other SSL methods, albeit in a non-federated setup. Table 2 compares the performance of FedGaLA with non-federated (centralized) methods for PACS and DomainNet datasets. As shown in this table, FedGaLA outperforms these methods by large margins. This is an expected observation as prior works have shown that federation can boost domain generalization [9, 54].
Model PACS DomainNet P A C S Ave. C I Q Ave. BYOL 27.0 25.9 21.0 19.7 23.4 14.6 8.7 5.9 9.7 MoCo 44.2 25.9 33.5 25.0 32.1 32.5 18.5 8.1 19.7 SimCLR 54.7 37.7 46.00 28.3 41.6 37.1 19.9 12.3 23.1 AdCo 46.5 30.2 31.5 22.9 32.8 32.3 17.9 11.6 20.6 DARLING 53.4 39.9 46.4 30.2 42.5 35.2 20.9 15.7 23.9 FedGaLA (ours) 64.7(1.9) 44.2(1.2) 60.5(2.2) 70.5(1.3) 60.0 47.6(0.9) 14.2(0.5) 61.4(0.3) 41.1
Models P A C S Ave.
@@
Figure 2: Covariance against domain shift for P, A, C, and S.
Gradient misalignment due to domain shift. In Figure 2 we illustrate the amount of measured covariance for different domains in PACS versus the amount of domain shift between each pair measured through Mutual Information. We observe that, except for a single outlier, the trend follows our prediction based on Theorem 1.
Analysis. All experiments below are performed on the PACS dataset using the domain ‘Photo’ unless otherwise mentioned. Additional analyses are also presented in Appendix E.
Ratio of labeled data. Figure 3(a) presents the results when evaluated with different ratios of labeled data, ranging from 10% to 60%. As can be seen, the accuracy increases significantly from 64% to 72% when the labeled ratio rises from 10% to 30%, followed by a steady climb to 74% as the labeled ratios increase to 60%. Overall, when more labeled data are used to train the linear classifier, the final performance expectedly improves.
@@
Figure 3: Impact of different labeled data ratios on linear evaluation (a) and average ratio of discarded local gradients over 100 communication rounds for $\tau=0$ and $E=1$ (b). The impact of the local threshold for different numbers of local epochs (c) and communication rounds (d) on performance.
Local threshold vs. the number of local epochs. In Figure 3(c) we study the impact of the threshold for the similarity between gradients and the reference ($\tau$) and the number of local epochs $E$ on performance. In this experiment, we use three different values for $\tau$: --0.1, 0, and 0.1, and three different values for $E$: 5, 7 or 9 (local epochs per communication round). We observe that our method produces the best results when $\tau=0$, i.e., when we keep all gradients with positive cosine similarity with the reference. Expectedly, even discarding gradients with small amounts of alignment ($\tau=0.1$) degrades the results, while keeping gradients that are not aligned with the reference ($\tau=-0.1$) also hurts performance. Moreover, we see that setting $E=7$ yields the best performance as increasing the number of epochs beyond 7 does not have a positive impact, and only increases computational time.
Communication rounds. We investigate the effect of the communication rounds $T$ on the model’s performance. Following [7, 27] we set $E=1$ and vary the number of communication rounds $T$ from 100 to 900. We observe from Figure 3(d) that performance improves significantly from 1 to 200 communication rounds, with this trend slowing down from 200 to 900.
Effects of regularizers on FedGaLA. Prior works have demonstrated that adding regularizers can indeed improved generalization across domains or non-IID data [16, 9]. To this end, we test the impact of regularizers on FedGaLA by applying two types of regularizers based on L2 norm [16] and FedProx [9]. Please refer to Appendix D.5 for more details regarding these two techniques. The results in Table 4 demonstrate that FedGaLA is highly compatible with regularizers and that the addition of such approaches can further boost the performance of our method.
@@
Table 4: The effect of regularizers on the performance of FedGaLA across PACS, DomainNet, Office-Home, and TerraInc datasets.
Model PACS DomainNet P A C S Ave. C I Q Ave. FedGaLA+L2 ($\lambda=0.001$) 64.8(0.5) 42.8(1.1) 58.7(1.6) 67.3(0.6) 58.4 46.3(0.8) 16.4(0.9) 62.8(0.2) 41.6 FedGaLa+L2 ($\lambda=0.01$) 66.5(0.9) 43.4(0.9) 56.9(2.7) 65.2(0.6) 57.9 38.9(1.2) 14.6(0.7) 52.5(0.8) 35.3 FedGaLA+FedProx ($\mu=0.001$) 62.1(0.2) 41.4(0.7) 58.9(3.1) 68.8(0.1) 57.8 44.6(0.2) 13.4(1.0) 60.9(1.1) 39.6 FedGaLA+FedProx ($\mu=0.01$) 63.1(0.2) 41.6(1.2) 58.5(2.0) 68.9(1.9) 58.0 44.7(1.2) 13.2(0.4) 61.7(0.4) 39.9 FedGaLA (ours) 64.7(1.9) 44.2(1.2) 60.5(2.2) 70.5(1.3) 60.0 47.6(0.9) 14.2(0.5) 61.4(0.3) 41.1
Model Office-Home TerraInc A C P R Ave. L38 L43 L46 L100 Ave. FedGaLA+L2 ($\lambda=0.001$) 11.8(0.4) 23.9(0.5) 37.1(0.4) 21.9(0.7) 23.6 61.6(0.9) 45.8(0.6) 46.6(0.7) 70.8(0.7) 56.2 FedGaLa+L2 ($\lambda=0.01$) 11.8(0.2) 21.2(0.1) 34.5(0.3) 22.1(0.2) 22.4 61.5(0.7) 38.1(2.9) 44.9(0.1) 57.8(3.3) 50.6 FedGaLA+FedProx ($\mu=0.001$) 10.2(0.4) 24.2(0.4) 36.4(1.1) 19.7(0.32) 22.6 63.8(0.7) 49.3(1.5) 47.2(0.1) 71.5(0.9) 57.9 FedGaLA+FedProx ($\mu=0.01$) 9.2(0.3) 23.6(0.7) 37.1(0.6) 20.4(0.4) 22.6 63.2(1.1) 47.8(2.1) 44.8(0.4) 71.4(0.8) 56.0 FedGaLA (ours) 8.9(0.4) 25.3(0.6) 36.6(0.3) 21.2(0.5) 23.0 63.6(0.1) 47.6(1.4) 43.9(2.9) 71.7(0.9) 56.7
FedGaLA with other SSL techniques. We examine the effectiveness of FedGaLA using other SSL frameworks (MoCo, BYOL, and SimCLR), on the PACS dataset and present the results in Appendix E. Our findings suggest that FedGaLA can improve the performance of various SSL baselines for federated unsupervised domain generalization.
@@ 6 Conclusion
Limitations. Given the emphasis of federated learning on privacy and safety, the robustness of FedGaLA against adversarial attacks, particularly those targeting the gradient exchange process, remains an unexplored area, which can be an interesting future research direction.
@@ A.1 Proof of Lemma 1
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\mathbb{E}\{(\textbf{g}_{i}-\mathbb{E}\{\textbf{g}_{i}\})(\textbf{g}_{j}-\mathbb{E}\{\textbf{g}_{j}\})\}.$ (1)
First, we calculate $\mathbb{E}{\textbf{g}_{i}}$. Given g is differentiable, as per the Assumption 1, we can estimate it using the first-degree Taylor expansion theorem around $\mu$:
$g(x)\approx g(\mu)+\sum_{f=1}^{F}(x^{f}-\mu^{f})\left(\frac{\partial g}{\partial x^{f}}\bigg{|}_{x^{f}=\mu^{f}}\right).$ (2)
We can use this equation to estimate $\mathbb{E}{\textbf{g}}$ as follows:
$\mathbb{E}{\textbf{g}}=\mathbb{E}{g(\mu)+\sum_{f=1}^{F}(x^{f}-\mu^{f})\left(\frac{\partial g}{\partial x^{f}}\bigg{|}_{x^{f}=\mu^{f}}\right)}.$ (3)
Note that $g(\mu)$ is constant, therefore $\mathbb{E}{g(\mu)}=g(\mu)$. Moreover, since we assume that the distribution of the features is normal, we can conclude:
$\mathbb{E}{\sum_{f=1}^{F}(x^{f}-\mu^{f})\left(\frac{\partial g}{\partial x^{f}}\bigg{|}_{x^{f}=\mu^{f}}\right)}=0.$ (4)
Since $\mathbb{E}{x^{f}-\mu^{f}}=0$ for $x^{f}\sim\mathcal{N}$, $\mathbb{E}{g}=g(\mu)$. By replacing $g_{i}$ and $g_{j}$ with their Taylor expansion we derive:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=$ $\mathbb{E}\left[\left(\sum_{f=1}^{F}(x_{i}^{f}-\mu_{i}^{f})\left(\frac{\partial g}{\partial x_{i}^{f}}\bigg{|}_{x_{i}^{f}=\mu_{i}^{f}}\right)\right)\right.\left.\left(\sum_{f=1}^{F}(x_{j}^{f}-\mu_{j}^{f})\left(\frac{\partial g}{\partial x_{j}^{f}}\bigg{|}_{x_{j}^{f}=\mu_{j}^{f}}\right)\right)\right].$ (5)
@@
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\mathbb{E}\left[\sum_{f=1}^{F}\sum_{e=1}^{F}(x_{i}^{f}-\mu_{i}^{f})(x_{j}^{e}-\mu_{j}^{e})\left(\frac{\partial g}{\partial x_{i}^{f}}\bigg{|}_{x_{i}^{f}=\mu_{i}^{f}}\right)\right.\left.\left(\frac{\partial g}{\partial x_{j}^{e}}\bigg{|}_{x_{j}^{e}=\mu_{j}^{e}}\right)\right].$ (6)
Given $\text{Cov}(x^{f}_{i},x^{e}_{j})=\mathbb{E}{(x^{f}_{i}-\mu^{f}_{i})(x^{e}_{j}-\mu^{e}_{j})}=\sigma_{x^{f}_{i},x^{f}_{j}}$, we derive:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\sum_{f=1}^{F}\sum_{e=1}^{F}\sigma_{x^{f}_{i},x^{f}_{j}}\left(\frac{\partial g}{\partial x_{i}^{f}}\bigg{|}_{x_{i}^{f}=\mu_{i}^{f}}\right)\left(\frac{\partial g}{\partial x_{j}^{e}}\bigg{|}_{x_{j}^{e}=\mu_{j}^{e}}\right).$ (7)
As the features are assumed to be independent, $\text{Cov}(x^{f}_{i},x^{e}_{j})=0$ when $f\neq e$, we have:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\sum_{f=1}^{F}\sigma_{x^{f}_{i},x^{f}_{j}}\left(\frac{\partial g}{\partial x_{i}^{f}}\bigg{|}_{x_{i}^{f}=\mu_{i}^{f}}\right)\left(\frac{\partial g}{\partial x_{j}^{f}}\bigg{|}_{x_{j}^{f}=\mu_{j}^{f}}\right).$ (8)
which completes the proof. ∎
A.2 Proof of Claim 1

### v2

@@ 5 Results
In this section, we first present the performance evaluation of FedGaLA across various datasets, comparing it against baseline methods to demonstrate its effectiveness. This is followed by ablation studies and sensitivity analyses, which provide detailed insights into the contributions of individual components and the robustness of our proposed framework.
5.1 Performance
Federated unsupervised domain generalization. We report the accuracy rates of FedGaLA and baseline models on the four datasets. As shown in Table 1, FedGaLA consistently outperforms all baselines across all four datasets, with the exception of the ‘Art-painting’ domain in Office-Home, for the 10% data regime. When 30% of the data are used, our method still generally outperforms the baseline models, although, for some of the domains, the baseline solutions produce slightly better results. This observation is expected given that with the introduction of more domain-specific training data, the need for domain generalization declines, and thus, methods that are not explicitly designed for domain generalization can produce competitive results. Across the four datasets, we observe that among the baseline models, FedSimCLR achieves better results compared to FedSimSiam, FedBYOL, and FedMoCo. This finding is consistent with (Zhang et al. 2022) where it was demonstrated that SimCLR provides a better foundation for domain generalization versus other SSL methods, albeit in a non-federated setup.
Centralized unsupervised domain generalization. We compare the performance of FedGaLA with centralized (non-federated) methods for PACS and DomainNet datasets, where the baselines are trained on the entire dataset consisting of all the domains. The results presented in Table 2 show that FedGaLA outperforms centralized methods by large margins. This is an expected observation as prior works have shown that federation can boost domain generalization (Nguyen, Torr, and Lim 2022; Arasteh et al. 2023).
Model P A C S Ave. FedMoCo 46.2(1.1) 28.6(1.0) 46.7(0.6) 37.6(1.3) 39.8 FedGaLA w/ MoCo 46.5(0.3) 28.9(0.4) 47.9(0.3) 39.6(2.4) 40.7 FedBYOL 52.1(1.1) 31.8(1.1) 45.4(2.2) 47.4(1.9) 44.2 FedGaLA w/ BYOL 52.8(0.6) 31.7(0.5) 46.3(3.1) 47.6(1.2) 44.6 FedSimCLR 64.2(1.2) 41.9(1.5) 58.4(1.3) 70.1(1.2) 58.6 FedGaLA w/ SimCLR 64.7(1.9) 44.2(1.2) 60.5(2.2) 70.5(1.3) 60.0
Model P A C S Ave.
@@ 5.1 Performance
Figure 3: (a) Impact of the number of iterations for global alignment. (b) Stability of the training. (c) Effect of different labeled data ratios on linear evaluation performance. (d) Performance of the model in different communication rounds. Results are reported for domain $P$ of PACS dataset with $10\%$ label ratio, except for (c) where the label ratio is changing.
FedGaLA with other SSL techniques. Recall that we used SimCLR in our federated framework. To further explore whether FedGaLA can improve the performance of other federated SSL methods, we perform additional experiments in which we employ MoCo and BYOL in FedGaLA instead of SimCLR. We evaluate the performance on the PACS dataset, as presented in Table 3. The results show that FedGaLA consistently improves the performance of all federated SSL frameworks, demonstrating its ability to adapt to different SSL methodologies while boosting domain generalization.
5.2 Empirical verification of Theorem 1
Gradient misalignment due to domain shift. In Figure 2(a) we illustrate the amount of measured covariance for different domains in PACS versus the amount of domain shift between each pair measured through Mutual Information. We observe that, except for a single outlier, the trend follows our prediction based on Theorem 1.
5.3 Ablation studies and component analysis
@@
Model Communication Frequency ($E$) $E=1$ $E=5$ $E=7$ $E=9$ FedGaLA 67.8 65.2 64.8 64.7
Table 5: Impact of communication frequency ($E$) on model accuracy over 900 total local epochs.
Models P A C S Ave.
RF: 0.001 64.3(1.9) 41.8(1.2) 58.1(2.3) 68.3(1.1) 58.1
RF: 0.01 63.8(1.4) 42.0(1.7) 58.1(0.4) 68.3(1.6) 58.6
RF: 0.1 64.1(1.7) 40.9(2.1) 58.5(1.4) 70.4(0.4) 58.5
FedGaLA (ours) 64.7(1.9) 44.2(1.2) 60.5(2.2) 70.5(1.3) 60.0
Table 6: Results on PACS dataset for different reweight-factors and our FedGaLA method. ‘RF’ indicates the re-weight factor.
Model PACS DomainNet
P A C S Ave. C I Q Ave.
FedGaLA+L2 ($\lambda=0.001$) 64.8(0.5) 42.8(1.1) 58.7(1.6) 67.3(0.6) 58.4 46.3(0.8) 16.4(0.9) 62.8(0.2) 41.6
FedGaLa+L2 ($\lambda=0.01$) 66.5(0.9) 43.4(0.9) 56.9(2.7) 65.2(0.6) 57.9 38.9(1.2) 14.6(0.7) 52.5(0.8) 35.3
FedGaLA+FedProx ($\mu=0.001$) 62.1(0.2) 41.4(0.7) 58.9(3.1) 68.8(0.1) 57.8 44.6(0.2) 13.4(1.0) 60.9(1.1) 39.6
FedGaLA+FedProx ($\mu=0.01$) 63.1(0.2) 41.6(1.2) 58.5(2.0) 68.9(1.9) 58.0 44.7(1.2) 13.2(0.4) 61.7(0.4) 39.9
FedGaLA (ours) 64.7(1.9) 44.2(1.2) 60.5(2.2) 70.5(1.3) 60.0 47.6(0.9) 14.2(0.5) 61.4(0.3) 41.1
Model Office-Home TerraInc
A C P R Ave. L38 L43 L46 L100 Ave.
FedGaLA+L2 ($\lambda=0.001$) 11.8(0.4) 23.9(0.5) 37.1(0.4) 21.9(0.7) 23.6 61.6(0.9) 45.8(0.6) 46.6(0.7) 70.8(0.7) 56.2
FedGaLa+L2 ($\lambda=0.01$) 11.8(0.2) 21.2(0.1) 34.5(0.3) 22.1(0.2) 22.4 61.5(0.7) 38.1(2.9) 44.9(0.1) 57.8(3.3) 50.6
FedGaLA+FedProx ($\mu=0.001$) 10.2(0.4) 24.2(0.4) 36.4(1.1) 19.7(0.32) 22.6 63.8(0.7) 49.3(1.5) 47.2(0.1) 71.5(0.9) 57.9
FedGaLA+FedProx ($\mu=0.01$) 9.2(0.3) 23.6(0.7) 37.1(0.6) 20.4(0.4) 22.6 63.2(1.1) 47.8(2.1) 44.8(0.4) 71.4(0.8) 56.0
@@
Table 7: The effect of regularizers on the performance of FedGaLA across PACS, DomainNet, Office-Home, and TerraInc.
5.4 Hyperparameter sensitivity and stability
Batch size. To identify the optimal training hyperparameters, we conduct a grid search on various batch sizes, testing values of 32, 64, 128, and 256, as illustrated in Figure 2(c). The results indicate that FedGaLA achieves its best performance with a batch size of 128, which balances efficient gradient updates and stability during training.
Local threshold vs. the number of local epochs. In Figure 2(d) we study the impact of the threshold for the similarity between gradients and the reference ($\tau$) and the number of local epochs $E$ on performance. In this experiment, we use three different values for $\tau$: --0.1, 0, and 0.1, and three different values for $E$: 5, 7 or 9 (local epochs per communication round). We observe that our method produces the best results when $\tau=0$, i.e., when we keep all gradients with positive cosine similarity with the reference. Expectedly, even discarding gradients with small amounts of alignment ($\tau=0.1$) degrades the results, while keeping gradients that are not aligned with the reference ($\tau=-0.1$) also hurts performance. Moreover, we see that setting $E=7$ yields the best performance as increasing the number of epochs beyond 7 does not have a positive impact, and only increases computational time.
Global gradient alignment iterations. Figure 3(a) shows the impact of different gradient alignment iterations (line 7 in Algorithm 1) on performance. FedGaLA produces the best results when the iteration number is 3. When the number of iterations increases, the model places too much emphasis on clients aligned with the global average and decreases the impact of unaligned clients harshly, leading to a decline in the overall performance of the model. On the other hand, with small iteration values, global gradient alignment is not performed effectively, and thus performance declines.
Training stability. Figure 3(b) demonstrates that FedGaLA maintains robust training stability, rapidly improving accuracy in the initial iterations and stabilizing after approximately 40 iterations. Despite discarding dissimilar gradients, the method avoids divergence and ensures consistent optimization, effectively learning domain-invariant features while maintaining steady performance throughout training.
5.5 Data and communication efficiency
Ratio of labeled data. Figure 3(c) presents the results when evaluated with different ratios of labeled data, ranging from 10% to 60%. As can be seen, the accuracy increases significantly from 64% to 72% when the label ratio rises from 10% to 30%, followed by a steady climb to 74% as the label ratios increase to 60%. Overall, when more labeled data are used to train the linear classifier, the final performance expectedly improves.
Communication rounds. We investigate the effect of the communication rounds $T$ on the model’s performance. Following (Zhuang et al. 2021; Zhuang, Wen, and Zhang 2022) we set $E=1$ and vary the number of communication rounds $T$ from 100 to 900. We observe from Figure 3(d) that performance improves significantly from 1 to 200 communication rounds, with this trend slowing down from 200 to 900.
FedGaLA and communication cost. We conduct an experiment to analyze the trade-off between communication frequency and model performance. In this setup, local models are trained for a total of 900 epochs, with aggregation occurring after every $E$ local epochs. As $E$ increases, communication frequency decreases, thereby reducing the overhead. The performance of the model is then assessed for different values of $E$, as demonstrated in Table 5. The results indicate that increasing the communication frequency (lower $E$) improves accuracy and the model maintains consistent performance even with reduced communication, particularly for $E>1$. This indicates the ability of the model to adapt to less frequent aggregation intervals.
5.6 Model variations
Gradient discarding vs. gradient re-weighting. Gradient re-weighting is an alternative approach to discarding unaligned gradients. To further highlight the effectiveness of gradient discarding in the local alignment module of FedGaLA, we conducted an additional experiment replacing gradient discarding with gradient re-weighting in FedGaLA. Here, rather than discarding the unaligned gradients, they are multiplied by a re-weight factor reducing their impact. The results are presented in Table 6, demonstrating that while re-weighting is effective, our original method of discarding unaligned gradients yields the best overall performance.
Effects of regularizers on FedGaLA. Prior works have demonstrated that adding regularizers can indeed improve generalization across domains or non-IID data (Li et al. 2020; Nguyen, Torr, and Lim 2022). To this end, we test the impact of regularizers on FedGaLA by applying two types of regularizers based on L2 norm (Li et al. 2020) and FedProx (Nguyen, Torr, and Lim 2022). Please refer to Appendix B.4 for more details regarding these two techniques. The results in Table 7 demonstrate that FedGaLA is highly compatible with regularizers and that the addition of such approaches can further boost the performance of our method.
@@ 6 Conclusion
Acknowledgments
We would like to thank Geotab Inc., the City of Kingston, and NSERC for their support of this work.
@@ A.1 Proof of Lemma 1
By definition, the Mutual Information between two random variables, $\textbf{x}_{i}$ and $\textbf{x}_{j}$ is
[…]
$p(\textbf{x}_{i},\textbf{x}_{j})=\prod_{f=1}^{F}p(x_{i}^{f},x_{j}^{f}).$ (7)
[…]
For bivariate Normal distributions $x_{i}$ and $x_{j}$, Mutual Information can be measured by $I(x_{i};x_{j})=-\frac{1}{2}\log(1-\rho^{2})$, where $\rho=\frac{\sigma_{x_{i},x_{j}}}{\sigma_{x_{i}}\sigma_{x_{j}}}$ is the correlation coefficient (Goldfeld and Greenewald 2021). Accordingly, Eq. (8) yields:
$I(\textbf{x}_{i};\textbf{x}_{j})=-\frac{1}{2}\sum_{f=1}^{F}\log(1-\sigma_{x_{i}^{f},x_{j}^{f}}^{2}),$ (10)
which completes the proof. ∎
A.2 Proof of Lemma 2
Proof.
@@ A.2 Proof of Lemma 2
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\mathbb{E}[(\textbf{g}_{i}-\mathbb{E}[\textbf{g}_{i}])(\textbf{g}_{j}-\mathbb{E}[\textbf{g}_{j}])^{T}].$ (11)
First, we calculate $\mathbb{E}[\textbf{g}_{i}]$. Given g is differentiable, as per the Assumption 1, we can estimate it using the first-degree Taylor expansion theorem around $\mu$:
$g(x)\approx g(\mu)+J_{g}(\mu).(x-\mu),$ (12)
where $J_{g}(.)$ is the Jacobian matrix $g$. We can use this equation to estimate $\mathbb{E}[\textbf{g}]$ as follows:
$\mathbb{E}[\textbf{g}]=\mathbb{E}\left[g(\mu)+J_{g}(\mu).(x-\mu))\right].$ (13)
Note that $g(\mu)$ is constant, therefore $\mathbb{E}[g(\mu)]=g(\mu)$. Moreover, since we assume that the distribution of the features is normal, we can conclude:
$\mathbb{E}\left[J_{g}(\mu).(x-\mu))\right]=\textbf{0}_{F}.$ (14)
Since $\mathbb{E}[x-\mu]=\textbf{0}_{F}$ for $x\sim\mathcal{N}$, $\mathbb{E}[g]=g(\mu)$. By replacing $\textbf{g}_{i}$ and $\textbf{g}_{j}$ with their Taylor expansion we derive:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\\ \mathbb{E}\left[\left(J_{g_{i}}(\mu_{i}).(x-\mu_{i})\right)\left.\right..\left(J_{g_{j}}(\mu_{j}).(x-\mu_{j}))\right)^{T}\right].$ (15)
@@
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=\\ J_{g_{i}}(\mu_{i}).\mathbb{E}\left[(x-\mu_{i}).(x-\mu_{j})^{T}\right].J_{g_{j}}(\mu_{j})^{T}.$ (16)
Given the covariance matrix $\Sigma=\mathbb{E}[(x-\mu_{i})(x-\mu_{j})]$, we derive:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})=$ $J_{g_{i}}(\mu_{i}).\Sigma.J_{g_{j}}(\mu_{j})^{T}.$ (17)
As the features are assumed to be independent, $\Sigma$ is a diagonal matrix, the general term of each entry of $\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})$ can be derived as:
$\text{Cov}(\textbf{g}_{i},\textbf{g}_{j})_{mn}=\\ \sum_{f=1}^{F}\sigma_{x^{f}_{i},x^{f}_{j}}\left(\frac{\partial g_{i(m)}}{\partial x_{i}^{f}}\bigg{|}_{\mu_{i}^{f}}\right)\left(\frac{\partial g_{j(n)}}{\partial x_{j}^{f}}\bigg{|}_{\mu_{j}^{f}}\right),$ (18)
where $g_{i(m)}$ and $g_{j(n)}$ represented the $m^{th}$ and the $n^{th}$ dimension of $g_{i}$ and $g_{j}$, respectively. ∎
A.3 Proof of Claim 1